OpenStudy (anonymous):

write an equation of the line perpendicular to the line given by 2x+3y=7 and having the same y -intercept as the line given by 5x+2y=10

OpenStudy (midhun.madhu1987):

The product of the slopes of 2 perpendicular lines is -1.

OpenStudy (midhun.madhu1987):

Find the slope of the line 2x + 3y = 7 To do so, convert this equation of the form y = mx + c where m is the slope

OpenStudy (anonymous):

so I would do 2x+3y=7 -2x -2x 3y=-2+7 then what would be my next step

OpenStudy (midhun.madhu1987):

 so what is the slope of this line?